### v1

@@ Abstract
Decoding nervous system activity is a key challenge in neuroscience and neural interfacing. In this study, we propose a novel neural decoding system that enables unprecedented large-scale sampling of muscle activity. Using micro-electrode arrays with more than 100 channels embedded within the forearm muscles, we recorded high-density signals that captured multi-unit motor neuron activity. This extensive sampling was complemented by advanced methods for neural decomposition, analysis, and classification, allowing us to accurately detect and interpret the spiking activity of spinal motor neurons that innervate hand muscles. We evaluated this system in two healthy participants, each implanted with three electromyogram (EMG) micro-electrode arrays (comprising 40 electrodes each) in the forearm. These arrays recorded muscle activity during both single- and multi-digit isometric contractions. For the first time under controlled conditions, we demonstrate that multi-digit tasks elicit unique patterns of motor neuron recruitment specific to each task, rather than employing combinations of recruitment patterns from single-digit tasks. This observation led us to hypothesize that hand tasks could be classified with high precision based on the decoded neural activity. We achieved perfect classification accuracy (100%) across 12 distinct single- and multi-digit tasks, and consistently high accuracy (>96%) across all conditions and subjects, for up to 16 task classes. These results significantly outperformed conventional EMG classification methods. The exceptional performance of this system paves the way for developing advanced neural interfaces based on invasive high-density EMG technology. This innovation could greatly enhance human-computer interaction and lead to substantial improvements in assistive technologies, offering new possibilities for restoring motor function in clinical applications.
@@
motor neuron, motor unit, intramuscular EMG, myoelectric control
@@ 1 Introduction
Neural interface research aims to restore or establish connections between the nervous system and the external environment, for example in individuals with neurological impairments [1, 2]. These interfaces, based on stimulating (encoding) and/or recording (decoding) from neural pathways, are used to assist individuals with disabilities or enhance the capabilities of healthy individuals. For this purpose, the decoding of neural signals provides commands to virtual systems, wearable robots or tele-operated devices.
@@
Direct brain signal decoding is the only option for severe impairments; however, for many other applications, nerve or muscle interfacing may be preferable [3, 4]. The surface electromyogram (EMG) is particularly simple to record and is used in several interfacing applications [5, 6, 7]. Although recording signals from muscle tissue is not a direct neural interface, the EMG captures the electrical activity produced by the muscles, which reflects the neural activation sent from the spinal cord’s output layers to the muscles [8]. Therefore, surface EMG signals can be decoded to reflect the activities of individual spinal motor neurons, which can then be mapped into estimated motor intent.
@@
Despite the potential of surface EMG-based decoding of neural activity, current decoding and mapping systems are limited in accuracy and, particularly, vary significantly in performance across individuals and conditions. The primary cause of this variability is the dependence of surface EMG characteristics on the volume conductor separating the sources of signal (muscle fibers) and the recording electrodes. Associated to this limitation are the sensitivity of surface EMG features to electrode misplacement, typically due to donning and doffing of the wearable device [9], and EMG signal cross-talk [10, 11]. These and other limitations of wearable EMG sensors can be addressed with invasive electrodes. Recent advances in intramuscular EMG technology have provided micro-electrode arrays with tens of electrodes closely spaced with respect to each other (high-density EMG micro-electrode arrays) [12, 13].
@@
Here, we present an innovative neural decoding system that extracts neural information from high-density intramuscular EMG (HD-iEMG) micro-electrode array signals, mapping precise discharge timings of the decoded motor neurons into the executed tasks. This system was experimentally validated offline on two healthy individuals, using three concurrently implanted intramuscular electrodes in each subject (Fig. 1). The validation involved classifying single- and multi-digit isometric tasks. The performance of the HD-iEMG mapping was then compared to the use of traditional pattern recognition techniques applied to the high-density surface EMG (HD-sEMG) and HD-iEMG signals. Therefore, we present four key advances in the field of neural decoding: 1) recordings of 120 intramuscular signals concurrently detected, paired with corresponding HD-sEMG signals. This is so far the largest reported number of implanted electrodes in the forearm muscles; 2) new perspectives on the neural command structure underlying finger movements; 3) a highly accurate neural classifier (100% accuracy in single- and multi-digit tasks for 12 classes), using physiological and interpretable features; 4) a comprehensive comparison of mapping performance across invasive and non-invasive high-density recording modalities. These findings enhance our understanding of neuromuscular control of the hand and have significant implications for neuroprosthetics and human-machine interfaces.
@@
Figure 1: Experimental setup: Three micro-electrode arrays were implanted into the forearm muscles of two participants. Three high-density surface electrode grids were placed on the skin overlying the intramuscular recording sites. Two load cells per finger were used to measure flexion and extension forces. Representative data are shown for intramuscular EMG, surface EMG, and force recordings. A schematic of the high-density intramuscular micro-electrode arrays, with a detailed zoom-in on the electrode tip, is shown on the bottom right.
@@ 2.1 Targeted muscle insertions
The recording of multiple high-density intramuscular data arrays represents one of the key innovations of this study. HD-iEMG enables recordings from a significantly larger number of sites compared to traditional needle or wire electrode arrays [14]. Furthermore, HD-iEMG signals exhibit greater spatial and temporal sparsity than HD-sEMG, allowing for more precise analysis of motor neuron activity [13]. In this study, three HD-iEMG electrodes were inserted in the forearm muscles of two healthy participants, with one electrode placed in each of the most relevant extrinsic muscles for finger flexion and extension. Interfacing with motor units that control hand function poses inherent challenges due to the precise targeting required for intramuscular electrode placement in the primary muscles responsible for dexterous control. Optimal electrode placement sites were determined via a magnetic resonance imaging (MRI)- and ultrasound-based anatomical assessment of the forearm. MRI was employed to gain a detailed understanding of the forearm muscle anatomy and to guide the insertion point and angle of the electrodes. Specific forearm points were tracked using MRI markers across scans, allowing precise localization of marker positions and detailed insights into forearm structure (Fig. 2 a). Each electrode array, comprising 40 platinum electrodes (140 $\mu m\times$ 40 $\mu m$) linearly distributed over 2 cm with 0.5 mm inter-electrode (IED) spacing [13], targeted muscles involved in finger movements. The insertion of each array was facilitated by a 25-gauge hypodermic needle, which was attached to the electrode with a guiding filament. Once the electrode was positioned, the filament was cut, allowing the needle to be removed while the electrode remained securely in the muscle.
@@
The first micro-electrode array targeted the extensor pollicis longus (EPL) and extensor digitorum communis (EDC) muscles, activated for the extension of the thumb and the digits respectively. A distal location in the forearm was chosen to allow concurrent monitoring of the two muscles. The second array was inserted transversely to target the extensor digiti minimi (EDM) and the extensor digitorum communis (EDC), employing an unconventional technique that represents a novel approach to intramuscular electrode placement, thereby maximizing coverage of the muscles relevant to digit extension. The third array targeted the flexor digitorum superficialis (FDS) muscle, chosen for its accessibility and essential role in flexion tasks.
@@
MRI provides a comprehensive anatomical view of the forearm, but practical constraints limit its use during needle insertions. Once the electrode insertion points and general positioning were identified using MRI, precise muscle targeting during the actual insertion became essential to avoid veins, arteries, and nerves. This necessitated real-time imaging guidance to ensure accurate and safe insertion. A portable ultrasound device (Butterfly iQ+) was used to precisely visualise electrode paths to target muscles in real-time (Fig. 2 c and d, for Subject 1 (S1) and Subject 2 (S2), respectively). Post-insertion, subjects performed maximal muscle contractions (MVC) as reference for submaximal tasks. This procedure also confirmed secure electrode placement and minimised the risk of displacement.
@@
While electrode placement followed the above rigorous procedure, some errors in the exact target placements occurred. For example, the limited field of view of the ultrasound probe impeded the full tracking of the micro-electrode array targeting the EDM for both subjects (panels c, d, second column). The extension of the insertion to the boundary of the Extensor Carpi Radialis Brevis (ECRB) suggests that the electrode might not have fully targeted the EDM, potentially failing to record specific information related to the little finger tasks. Moreover, although the EPL was targeted in both subjects (panels c, d, first column), the number of channels within the muscle was relatively small (panels c, d first column). This potentially impacted the amount of information recorded from the EPL during these tasks. The micro-electrode arrays were acutely implanted in the target muscles with percutaneous wires.
Figure 2: MRI and ultrasound images of the forearm with the boundaries of the relevant muscles highlighted in color (EPL: pink, EDC: blue, EDM: orange, ECRB: green, FDS: red). a MRI scans of the forearm of one subject with highlighted boundaries of the targeted muscles. The first MRI image includes a marker for precise positioning within the forearm’s anatomical structure. b Detailed MRI view of the forearm. c, d Ultrasound image during the micro-electrode array insertion for subject S1 (c) and subject S2 (d). The guiding needle is highlighted in white.
@@ 2.2 Data characteristics
Concurrent recordings of HD-iEMG and HD-sEMG were obtained during flexion and extension tasks for each finger of the hand individually and for four combinations of finger tasks. By accurately decoding the electrical activity into the spiking activities of spinal motor neurons, we uncovered distinct neural signatures associated with each task, providing deeper insights into the neural mechanisms underlying fine motor control. Leveraging this new understanding, we then performed classification analyses across multiple conditions, further validating the robustness of these task-specific neural patterns. First, we classified 16 classes of gestures, which included all individual finger movements and three two-finger combinations, in both flexion and extension. Next, we analysed the 12 classes that corresponded to tasks primarily controlled by the implanted muscles, ensuring that these tasks involved movements whose neural activity was captured by the intramuscular micro-electrode arrays. Finally, as a proof of concept for prosthetic hand control (one representative application of this mapping system), we selected eight functionally relevant tasks, which are challenging to decode with current methods (see Section 4.1.2 in Methods for more details).
@@
Figure 3 a shows a representative example of the data recorded from one of the micro-electrode arrays within the muscle. The multi-channel data recorded from all three micro-electrode arrays was subsequently decomposed to identify the spiking times of individual motor units (Fig. 3 b). On average, 27.1±12.9 (S1) and 8.0±6.1 (S2) motor units were identified from the HD-iEMG per task (mean and std across two repetitions, the three electrodes, and the 16 tasks). The average discharge rate of the decomposed motor units across the two participants was 11.0±3.9 spikes/s. Figure 3 c presents the average amplitude maps of motor unit activation during different finger extension tasks. These maps were generated by first computing the spike-triggered average for each motor unit action potential, per electrode, and per task, followed by normalisation and rectification. The spike-triggered averages were then averaged across all active motor units for each electrode and task to highlight the areas of the electrode array that were most active for each task. As shown in Figure 3 c, the electrodes exhibited task-specific activity, indicating localised activation within each array. This observation reflects the high spatial selectivity of each HD-iEMG channel. While some improvements in insertion procedures may be needed in future studies, the results indicated that the recorded information was of high quality for the subsequent analyses.
@@
Figure 3: a A 0.5-second segment of HD-iEMG data captured during the extension of the little finger from the EPL electrode. b Raster plots showing the extension movements of individual fingers for the second repetition for subject S1. The sections within the dashed lines indicate the parts of the signal considered for this study. c Average amplitude map of motor unit activities during the extension task of each finger and for each electrode (S1). The x-axis represents time in ms, and the y-axis represents the 40 channels recorded per electrode.
2.3 Neural control of digit tasks
This study marks the first instance in which we have been able to sample a relatively large number of motor neurons across multiple single- and multi-digit tasks under standardized conditions. By analyzing the activities of these motor neurons, we uncovered distinct neural patterns associated with different tasks, enhancing our understanding of the neural mechanisms underlying fine motor control.
Figure 4 shows the distribution of active motor units for individual finger movements and three two-finger combinations, with flexion tasks in panel a and extension tasks in panel b. Tasks are arranged along the axes, with bold squares highlighting motor units engaged during the specific task. The data reveal that the highest number of active motor units were consistently task-specific. Motor units active during other tasks are shown using distinct colors and patterns—solid colors for flexion and diagonal hatching for extension—facilitating comparisons of motor unit engagement across different movements. While some motor units were engaged during other tasks, their contribution was minor compared to the dominant task-specific activity, emphasizing the precision and task-specific recruitment of motor units. Notably, motor units involved in single-digit tasks were rarely recruited for multi-digit tasks involving the same finger. On average, motor units from single-digit tasks contributed for only 2.3$\pm$3.7of the motor units recruited across all multi-digit tasks and movement directions, underscoring the specialized recruitment of motor units for multi-digit movements. Furthermore, motor units that were active across all tasks, but not specifically tied to the task in focus, contributed for 44.6$\pm$11%, though task-specific motor unit activity consistently accounted for the majority of the recruitment. These findings highlight the nuanced and highly selective recruitment patterns of motor units during complex hand movements.
Figure 4 c provides a representative example of motor unit action potential waveforms for index flexion (blue), middle flexion (green), and the combination of index and middle finger flexion (pink). The distinct waveform shapes between the combination and individual finger movements visually confirm, for this representative example, that mostly unique motor units were activated for each task. This suggests that specific motor units are recruited in different patterns depending on the complexity and nature of the finger movements. The ability to distinctly identify these patterns is crucial for accurate decoding and mapping systems. The distinct activation of motor units for each task suggests that this decoding system is well-suited for highly accurate gesture classification. This consistency supports the use of motor unit activity to reliably differentiate gestures, as explored in the subsequent classification analysis.
Figure 4: Motor unit counts for S1 during flexion (a) and extension (b) tasks classified from the 16 classes. Each bar represents the number of motor units that were active during the specific task, labelled according to the filter that produced the accepted source. The hatch pattern distinguishes the extension task from the flexion task, providing a clear visual differentiation between the two movements. This highlights the distribution and activity levels of motor units across different tasks. c Representative motor unit action potentials decomposed from the FDS electrode for S1 during various tasks: index finger flexion (blue), middle finger flexion (green), and the combination of index and middle finger flexion (pink). The majority of the motor units in the combination task originate from unique MUs specific to the combined movement, rather than being a combination of the motor units active during the individual index and middle finger flexion tasks. This indicates distinct motor unit recruitment patterns for the combined task compared to the individual tasks.

### v2

@@ Abstract
Decoding the activity of the nervous system is a critical challenge in neuroscience and neural interfacing. In this study, we present a neuromuscular recording system that enables large-scale sampling of muscle activity using microelectrode arrays with over 100 channels embedded in forearm muscles. These arrays captured intramuscular high-density signals that were decoded into patterns of activation of spinal motoneurons. In two healthy participants, we recorded high-density intramuscular activity during single- and multi-digit contractions, revealing distinct motoneuron recruitment patterns specific to each task. Based on these patterns, we achieved perfect classification accuracy (100%) for 12 single- and multi-digit tasks and over 96% accuracy for up to 16 tasks, significantly outperforming state-of-the-art EMG classification methods. This intramuscular high-density system and classification method represent an advancement in neural interfacing, with the potential to improve human-computer interaction and the control of assistive technologies, particularly for replacing or restoring impaired motor function.
@@
motoneuron, motor control, intramuscular EMG, human-machine interfaces
@@ 1 Introduction
Neural interface research aims to restore or establish connections between the nervous system and the external environment, for example, in individuals with neurological impairments (1, 2). These interfaces, based on stimulating (encoding) and/or recording (decoding) of neural pathways, are used to assist individuals with disabilities or enhance the capabilities of healthy individuals. For this purpose, the decoding of neural signals provides commands to virtual systems, wearable robots, or tele-operated devices.
@@
Direct brain signal decoding is the only option for severe impairments; however, for many other applications, nerve or muscle interfacing may be preferable (3, 4). The surface electromyogram (EMG) is particularly simple to record and is used in several interfacing applications (5, 6, 7). Although recording signals from muscle tissue is not a direct neural interface, the EMG captures the electrical activity produced by muscles, which reflects the neural activation sent from the spinal cord’s output layers to the muscles (8). Therefore, surface EMG signals can be decoded to reflect the activities of individual spinal motoneurons, which can then be mapped into estimated motor intent.
@@
Despite the potential of surface EMG-based decoding of neural activity, current decoding and mapping systems are limited in accuracy and, particularly, vary significantly in performance across individuals and conditions. The primary cause of this variability is the dependence of surface EMG characteristics on the volume conductor separating the sources of signal (muscle fibers) and the recording electrodes. Associated with this limitation are the sensitivity of surface EMG features to electrode misplacement, typically due to donning and doffing of the wearable device (9), and EMG signal cross-talk (10, 11). These and other limitations of wearable EMG sensors can be overcome with invasive electrodes. Recent advances in intramuscular EMG technology have provided microelectrode arrays with tens of electrodes closely spaced with respect to each other (high-density EMG microelectrode arrays) (12, 13).
@@
Here, we present an innovative neural decoding system that extracts neural information from high-density intramuscular EMG (HD-iEMG) microelectrode array signals, mapping precise discharge timings of the decoded motoneurons into the executed tasks. This system was experimentally validated offline in two healthy individuals, using three concurrently implanted intramuscular electrodes in each participant (Fig. 1). The validation involved classifying single- and multi-digit isometric tasks. The performance of HD-iEMG mapping was then compared to the use of traditional pattern recognition techniques applied to high-density surface EMG (HD-sEMG) and HD-iEMG signals. Therefore, we present three key advances in the field of neural interfacing: 1) recordings of 120 intramuscular signals concurrently detected, paired with corresponding HD-sEMG signals. This is so far the largest reported number of implanted electrodes in the forearm muscles; 2) a highly accurate neural classifier (100% accuracy in single- and multi-digit tasks for 12 classes), using physiological and interpretable features; 3) a comprehensive comparison of mapping performance across invasive and non-invasive high-density recording modalities. These findings enhance our understanding of neuromuscular control of the hand and have significant implications for neuroprosthetics and human-machine interfaces.
@@
Figure 1: Experimental setup and intramuscular high-density EMG recordings. Three microelectrode arrays were implanted into the forearm muscles of two participants and three high-density surface electrode grids were placed on the skin overlying the intramuscular recording sites. Two load cells per finger were used to measure flexion and extension forces. Representative data are shown for intramuscular EMG, surface EMG, and force recordings. A schematic of the high-density intramuscular microelectrode arrays, with a detailed zoom-in on the electrode tip, is shown on the bottom right.
@@ 2.1 Targeted muscle insertions
The recording of multiple high-density intramuscular microelectrode arrays represents one of the key innovations of this study. HD-iEMG enables recordings from a significantly larger number of sites compared to traditional needle or wire electrode arrays (14). Furthermore, HD-iEMG signals exhibit greater spatial and temporal sparsity than HD-sEMG, allowing for more precise analysis of motoneuron activity (13). In this study, three HD-iEMG electrodes were inserted into the forearm muscles of two healthy participants, with one electrode placed in each of the most relevant extrinsic muscles for finger flexion and extension. While several other forearm muscles contribute to these movements, each electrode insertion is technically challenging and time-consuming; therefore, we selected three key muscles to balance feasibility with exhaustive signal acquisition. Interfacing with motoneurons that control hand function poses inherent challenges due to the precise targeting required for intramuscular electrode placement in the muscles responsible for dexterous control. Optimal electrode placement sites were determined via a magnetic resonance imaging (MRI)- and ultrasound-based anatomical assessment of the forearm. MRI was employed to gain a detailed understanding of the forearm muscle anatomy and to guide the insertion point and angle of the needles. Specific forearm points, based on anatomical landmarks, were tracked using MRI markers across scans, allowing precise localisation of marker positions and detailed insights into forearm structure (Fig. 2 A, B). Identification of the muscles with the ultrasound was aided by having participants performing voluntary contractions of the targeted muscles. Each electrode array, comprising 40 platinum electrodes (140 $\mu m$ $\times$ 40 $\mu m$) linearly distributed over 2 cm with 0.5 mm inter-electrode (IED) spacing (13), targeted muscles involved in finger flexion and extension. The insertion of each array was facilitated by a 25-gauge hypodermic needle, which was attached to the electrode with a guiding filament. Once the electrode was positioned, the filament was cut, allowing the needle to be removed while the electrode remained securely in the muscle (13).
@@
The first microelectrode array targeted the extensor pollicis longus (EPL) and extensor digitorum communis (EDC) muscles, activated for the extension of the thumb and the digits, respectively. A distal location in the forearm was chosen to allow concurrent monitoring of the two muscles with the same microelectrode. The second array was inserted transversely to target the extensor digiti minimi (EDM) and the extensor digitorum communis (EDC), employing an unconventional technique that maximises coverage of the muscles relevant to digit extension. The third array targeted the flexor digitorum superficialis (FDS) muscle, chosen for its accessibility and essential role in flexion tasks.
@@
MRI provides a comprehensive anatomical view of the forearm, but practical constraints limit its use during needle insertions. Once the electrode insertion points and general positioning were identified using MRI, precise muscle targeting during the actual insertion was critical to avoid veins, arteries, and nerves. This necessitated real-time imaging guidance to ensure accurate and safe insertion. A portable ultrasound device (Butterfly iQ+) was used to visualise the electrode paths to the target muscles in real-time (Fig. 2 C and D, for Participant 1 and Participant 2, respectively). Since the EDC and FDS have multiple compartments supplying different fingers, it is difficult to determine the exact location being sampled; however, we attempted to target all compartments by instructing participants to perform voluntary movements and verifying muscle activation. Post-insertion, participants performed maximal muscle contractions (MVC) as a reference for submaximal tasks. This procedure also confirmed secure electrode placement and minimised the risk of displacement.
@@
While electrode placement followed the above rigorous procedure, some errors in the exact target placements occurred. For example, the limited field of view of the ultrasound probe impeded the full tracking of the microelectrode array targeting the EDM for both participants (Fig. 2 C, D, second row). The extension of the insertion to the boundary of the Extensor Carpi Radialis Brevis (ECRB) suggests that the electrode might not have fully targeted the EDM, potentially failing to record specific information related to the little finger tasks. Moreover, although the EPL was targeted in both participants (Fig. 2 C, D, first row), the number of channels within the muscle was relatively small. This potentially impacted the amount of information recorded from the EPL during these tasks.
Figure 2: Anatomical assessment of forearm muscles and ultrasound-guided insertions. Muscle boundaries are highlighted in colour (EPL: pink, EDC: blue, EDM: yellow, ECRB: green, FDS: red). A Posterior MRI of the right forearm. Yellow lines mark the levels of each cross-sectional scan in panel B. B Forearm MRI scans from one participant with targeted muscles outlined in colour. The first and second columns include a marker for tracking anatomical landmarks across views. The ulna and radius indicate the forearm level at which each cross-sectional slice was obtained. C, D Ultrasound images during microelectrode array insertion for participant one (C) and participant two (D). The guiding needle is highlighted in white.
@@ 2.2 Data characteristics
Concurrent recordings of HD-iEMG and HD-sEMG were obtained during flexion and extension tasks for each finger of the hand individually and for four combinations of finger tasks. Two repetitions were recorded for each task. By decoding the electrical activity into the spiking activities of spinal motoneurons, we performed classification analyses across multiple conditions. First, we classified 16 classes of tasks, which included all individual finger contractions and three two-finger combinations, in both flexion and extension. Next, we analysed the 12 classes that corresponded to tasks primarily controlled by the implanted muscles, ensuring that these tasks involved contractions whose neural activity was captured by the intramuscular microelectrode arrays. Finally, as a proof of concept for prosthetic hand control (one representative application of this mapping system), we selected eight functionally relevant tasks, which are challenging to decode with current methods (see Section 4.1.2 in Methods for more details).
@@
Figure 3A presents a representative example of signals recorded from one of the intramuscular microelectrode arrays during two different extension tasks, illustrating their spatial distribution across the array. The multi-channel data recorded from all three microelectrode arrays was independently decomposed to identify the spiking times of individual motoneurons (Fig. 4A, 3B). On average, 27.1±12.9 (participant one) and 8.0±6.1 (participant two) motoneurons were identified from the HD-iEMG per task (mean and std across two repetitions, the three electrodes, and the 16 tasks). The average discharge rate of the decomposed motoneurons across the two participants was 11.0±3.9 pulses/s. In contrast, the HD-sEMG recordings yielded 5.1±4.4 (participant one) and 1.8±1.8 (participant two) motoneurons on average. The average discharge rate of the decomposed motoneurons across the two participants was 8.8±3.1 pulses/s. Due to the low number of motoneurons identified from the surface recordings, the subsequent analysis was limited to motoneurons identified from the intramuscular recordings. Figure 3C presents the summed amplitude maps of motoneuron activation during different finger extension tasks. These maps were generated by first computing the spike-triggered average for each motoneuron action potential, per electrode, and per task, followed by normalisation and rectification. The spike-triggered averages were then summed across all active motoneurons for each electrode and task to highlight the areas of the electrode array that were most active for each task. As shown in Figure 3C, the electrodes exhibited task-specific activity, indicating localised activation within each array. This observation reflects the high spatial selectivity of each HD-iEMG channel. Notably, the EPL exhibited only slight activity during thumb extension but was more active during the extension of the other digits. This pattern aligns with expectations (see Fig. 2), as the EPL electrode placement did not optimally capture the targeted muscle. In contrast, the EDC was selectively active during the extension of all digits except the thumb, consistent with its anatomical function as the primary extensor of the index, middle, ring, and little fingers. Additionally, substantial co-contraction of the flexor muscles was observed during digit extension, a phenomenon that is expected to stabilise the wrist and fingers during the contraction. This was particularly evident in the FDS, which exhibited higher activation, likely due to its large muscle volume and functional role in flexion and stabilisation. While some improvements in insertion procedures may be needed in future studies, the results indicated that the recorded information was of high quality for the subsequent analyses.
@@
Figure 3: Data characteristics. A A 0.5-second segment of HD-iEMG data captured during the extension from the EPL electrode of the little finger (left) and of the ring finger (right). B Raster plots showing the neural activity extracted from the extension contractions of individual fingers for the second repetition for participant one. The sections within the dashed lines indicate the parts of the signal considered for this study. The y-axis for the motoneurons is not shared across plots. C Summed amplitude map of motoneuron activities during the extension task of each finger and for each electrode (participant one).